There are 5 blue marbles and 7 orange marbles in a jar. Misha draws one blue marble and sets it aside. What is the probability that Misha will randomly draw an orange marble from the remaining marbles in the jar?

1
Expert's answer
2021-05-18T17:53:12-0400

Total number of marble"=5+7=12"


Since one blue ball is drawn. So remaining ball in jar="11"


Probability that Misha draw orange ball after "=\dfrac{7}{11}=0.636"